OpenStudy (anonymous):

which of the following is an x-intercept of the function, f(x)=x^2-25 A. -20 B. -15 C. -25 D. -5

OpenStudy (welshfella):

x^2 - 25 = 0 solve for x

OpenStudy (welshfella):

x intercepts occur when f(x) = 0

OpenStudy (anonymous):

-5

OpenStudy (welshfella):

x^2 - 25 = 0 add 25 to both sides x^2 = 25 x = ? can you complete this?

OpenStudy (welshfella):

one of the options fit this

OpenStudy (anonymous):

-5?

OpenStudy (welshfella):

yes beacuse -5 * -5 = 25

OpenStudy (anonymous):

yeahh I got it.. Thank you :p

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

yeahh I got it.. Thank you :p